OpenStudy (anonymous):

x^2-6x-3=4? How do I solve by completing the square?

OpenStudy (anonymous):

add 3 to both sides first

OpenStudy (anonymous):

get \[x^2-6x=7\] then take half of 6, which is 3, square it to get 9 and add it to both sides giving \[(x-3)^2=7+9=16\] then take the square root of both sides, don't forget the \(\pm\) etc

OpenStudy (anonymous):

Thanks! Just one question though, when I add 3 to both sides wouldn't i get =1 instead of =7?